OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. slope 4, and passes through (2, 20)

OpenStudy (anonymous):

y=mx+b, mis the slope and b is the y-intercept

OpenStudy (anonymous):

substitute y and x into y=mx+b to find the value of b

OpenStudy (anonymous):

put in all the values that you have been given: m=4 x=2 y=20 y=mx+b 20=2(4)+b 20=8+b 12=b y=4x+12